OpenStudy (anonymous):

If x=3 and y=4 then what is the value of r?

OpenStudy (lgbasallote):

do you have an equation?

OpenStudy (lgbasallote):

unless you mean \[r = \sqrt{x^2 + y^2}\]

OpenStudy (unklerhaukus):

\[r^2=x^2+y^2=3^2+4^2\]

OpenStudy (anonymous):

my answer choices are 25 SR7 3SR3 5 SR5

OpenStudy (unklerhaukus):

one of them is right